### v3

# EasyRec: Simple yet Effective Language Models for Recommendation

Abstract.
Deep neural networks have become a powerful technique for learning representations from user-item interaction data in collaborative filtering (CF) for recommender systems. However, many existing methods heavily rely on unique user and item IDs, which limits their ability to perform well in practical zero-shot learning scenarios where sufficient training data may be unavailable. Inspired by the success of language models (LMs) and their strong generalization capabilities, a crucial question arises: How can we harness the potential of language models to empower recommender systems and elevate its generalization capabilities to new heights? In this study, we propose EasyRec - an effective and easy-to-use approach that seamlessly integrates text-based semantic understanding with collaborative signals. EasyRec employs a text-behavior alignment framework, which combines contrastive learning with collaborative language model tuning, to ensure a strong alignment between the text-enhanced semantic space and the collaborative behavior information. Extensive empirical evaluations across diverse real-world datasets demonstrate the superior performance of EasyRec compared to state-of-the-art alternative models, particularly in the challenging text-based zero-shot recommendation scenarios. Furthermore, the study highlights the potential of seamlessly integrating EasyRec as a plug-and-play component into text-enhanced collaborative filtering frameworks, thereby empowering existing recommender systems to elevate their recommendation performance and adapt to the evolving user preferences in dynamic environments. For better result reproducibility of our EasyRec framework, the model implementation details, source code, and datasets are available at the link: https://github.com/HKUDS/EasyRec.
copyright: none
1. Introduction
Deep learning has established itself as a highly promising and powerful solution for capturing user preferences in the context of online recommender systems (Yang et al., 2022; Zhang et al., 2022). This approach harnesses the power of deep neural networks to learn rich and meaningful user and item representations by analyzing the complex patterns of user-item interactions. This, in turn, enables recommendation algorithms to accurately infer user preferences and provide highly relevant and personalized recommendations (Xu et al., 2023; Sun et al., 2021).
In recent years, there has been a notable surge of advancements in enhancing recommender systems through the incorporation of neural network-powered collaborative filtering frameworks, particularly in the domain of graph neural networks (GNNs) (Wang et al., 2019; He et al., 2020; Xia et al., 2022). By effectively leveraging the inherent graph structure present in the data, GNNs have demonstrated exceptional capabilities in capturing high-order relationships and complex dependencies among users and items. Notable examples of such GNN-based approaches include NGCF (Wang et al., 2019) and LightGCN (He et al., 2020). These methods have showcased impressive performance in recommendation tasks by virtue of their ability to model the collaborative signals present in the data through recursive message passing mechanisms.
The issue of data scarcity in recommender systems poses a significant challenge for existing deep collaborative filtering models, hindering their ability to learn precise user/item representations, particularly when dealing with sparse interaction data (Lin et al., 2021; Wei et al., 2021; Ren et al., 2024b; Hao et al., 2021). This challenge primarily arises from the limited availability of training labels, which in turn impedes the models’ capacity to capture the intricate relationships and dependencies that exist between users and items. To alleviate data scarcity effects, recent studies explored the potential of self-supervised learning to provide effective data augmentation. For example, in terms of contrastive augmentation, methods like SGL (Wu et al., 2021) and NCL (Lin et al., 2022) leverage graph contrastive learning to supplement the supervised recommendation task with the cross-view mutual information maximization. For generative augmentation, approaches such as AutoCF (Xia et al., 2023) are built upon the masked autoencoding mechanism, enabling the models to reconstruct the interaction structures for self-supervision.
While the recent advancements in self-supervised learning techniques have offered promising avenues for mitigating the impact of data scarcity in current collaborative filtering models, they also come with a significant limitation (Yuan et al., 2023). This limitation stems from the inherent design of the models, which heavily relies on using unique identities (IDs) to represent users and items throughout the entire process of learning representations. In practical recommenders, however, we often come across new recommendation data collected from different domains or time periods, involving diverse sets of users and items. This creates a challenge as existing ID-based recommendation models struggle to effectively incorporate and adapt to such new data, especially when there is a change in the set of identity tokens of users and items for zero-shot recommendation. The rigid dependence on user and item IDs in these models hinders their ability to generalize and perform well in scenarios where the user and item spaces are not static or fully overlapping.
For instance, the constant generation of new items like videos and social media content necessitates accurate recommendations in real-life recommender systems, even when there are limited interaction observations. While cross-domain recommendation methods (Xie et al., 2022; Cao et al., 2022) draw inspiration from leveraging information and knowledge across multiple domains to enhance recommendation performance, they often assume that users from different domains belong to the same set (Dacrema et al., 2012; Xie et al., 2022). Unfortunately, this assumption significantly restricts the flexibility and generalization capabilities of recommenders. As a result, existing methods may encounter difficulties in adapting and delivering accurate recommendations when confronted with diverse user populations across different domains.
Language Models as Zero-Shot Recommenders. Considering the challenges and motivations discussed earlier, the objective of this study is to introduce a recommender system that functions as a zero-shot learner, possessing robust generalization capabilities and the flexibility to adapt to new recommendation data seamlessly. To accomplish this objective, we propose integrating language models with collaborative relation modeling, forming an effective text embedder-EasyRec that is both lightweight and effective. This integration seamlessly combines text-based semantic encoding with high-order collaborative signals, resulting in a recommender system that offers a strong model generalization ability.
Recent research has explored leveraging large language models (LLMs) to enhance recommender systems. Existing approaches broadly fall into two categories. The first uses LLMs for data augmentation (e.g., RLMRec (Ren et al., 2024a), AlterRec (Li et al., 2024)), encoding textual information to complement collaborative filtering. While this combines LLM and collaborative strengths, these methods remain ID-based and struggle to generalize. The second approach utilizes LLMs to directly generate user-item interaction predictions based on language (e.g., LLaRA (Liao et al., 2024), CoLLM (Zhang et al., 2023)). However, such LLM-based recommenders face significant real-world limitations, suffering from poor efficiency (e.g., $\sim$1 second per prediction), rendering them impractical for large-scale recommendation tasks. These challenges highlight the need for more efficient, scalable solutions that seamlessly integrate the semantic understanding of textual information with the collaborative strengths of zero-shot recommenders.
Our proposed model has demonstrated superior performance in comparison to state-of-the-art language models, as illustrated in Figure 1. This performance advantage is achieved within a cost-efficient parameter space, ranging from 100 to 400 million parameters, with the computational cost of $\sim$0.01 second per prediction. Notably, our model exhibits the scaling law phenomenon, where its performance continually improves as the parameter size is increased. Furthermore, in contrast to existing approaches that suffer from poor efficiency, our model is designed to be highly scalable and practical for large-scale recommendation tasks.
Figure 1. EasyRecdemonstrates superior performance compared to state-of-the-art language models in text-based zero-shot recommendation on the Amazon and Steam datasets.
Figure 2. The overall framework of our proposed collaborative information-guided language model EasyRec.
In summary, this work makes the following key contributions:
  • Motivation. The primary objective of this study is to introduce a novel recommender system that is built upon language models to function as a zero-shot learner. This innovative approach aims to demonstrate exceptional adaptability to diverse recommendation data, while also exhibiting robust generalization capabilities.
  • Methodology. To align text-based semantic encoding with collaborative signals from user behavior, we propose a novel contrastive learning-powered collaborative language modeling approach, which allows the system to capture both the semantic representations of users/items, as well as the underlying behavioral patterns and interactions within the recommendation data.
  • Zero-Shot Recommendation Capacity. The EasyRec is extensively evaluated through rigorous experiments as a text-based zero-shot recommender. Performance comparisons reveal that it consistently achieves significant advantages over baseline methods in terms of recommendation accuracy and generalization capabilities. Furthermore, the study demonstrates the model’s remarkable potential in generating dynamic user profiles that are highly adaptive to time-evolving user preferences.
  • Existing CF Model Enhancement. Our proposed EasyRec has been seamlessly integrated as a lightweight, plug-and-play component with state-of-the-art collaborative filtering (CF) models. The lightweight and modular design of our language model is a key strength, as it facilitates the adoption of our novel recommendation paradigm across a wide range of use cases.
2. Preliminaries
In recommender systems, we have a set of users $\mathcal{U}$ and a set of items $\mathcal{I}$, along with the interactions between them (e.g., clicks, purchases). For each user $u\in\mathcal{U}$, we define $\mathcal{N}_{u}$ as the set of items that user $u$ has interacted with. Likewise, for each item $i\in\mathcal{I}$, we define $\mathcal{N}_{i}$ as the set of users who have interacted with that item. To represent these user-item interactions, we can use an interaction matrix $\mathcal{A}^{|\mathcal{U}|\times|\mathcal{I}|}$, where the entry $\mathcal{A}^{u,i}$ is 1 if user $u$ has interacted with item $i$, and 0 otherwise. The primary goal of a recommender model is to estimate the probability $p_{u,i}$ of a future interaction between a user $u$ and an item $i$. This predicted probability can then be used to generate personalized item recommendations for each user, tailored to their individual preferences and past behavior.
Text-based Zero-Shot Recommendation is essential in recommender systems, as it can address the common cold-start problem. In real-world scenarios, new users and new items often lack sufficient interaction data, making it challenging to provide effective personalized recommendations. By leveraging the textual descriptions of users and items, such as product titles, details, and user profiles, the language models can construct semantic representations to enable text-based recommendations for these cold-start situations. This overcomes the limitations of traditional collaborative filtering methods, by offering a distinct advantage over the traditional ID-based paradigm. By leveraging robust language models, this approach exhibits remarkable potential in “zero-shot” scenarios, where the testing data have not been previously encountered.
Formally, we define $\mathcal{P}_{u}$ and $\mathcal{P}_{i}$ as the generated text-based profiles of user $u$ and item $i$, respectively, which are encoded into representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ using a language model, as shown as:
(1) $\mathbf{e}_{u}=\text{LM}(\mathcal{P}_{u}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{e}_{i}=\text{LM}(\mathcal{P}_{i}).$
The interaction preference score $p_{u,i}$ between a user $u$ and an item $i$ can be calculated as the cosine similarity between their text embeddings $\mathbf{e}_{u}$ and $\mathbf{e}i$, i.e., $p_{u,i}=cos(\mathbf{e}_{u},\mathbf{e}_{i})$. We can then recommend to the user the top-$k$ uninteracted items with the highest similarity scores, yielding a personalized recommendation set.
(2) $\mathcal{R}_{u}=\text{top\text{-}k}_{i\in\mathcal{I}\setminus\mathcal{N}_{u}}cos(\mathbf{e}_{u},\mathbf{e}_{i}).$
Text-enhanced Collaborative Filtering. Collaborative filtering (CF) is a widely used recommender system approach that leverages the collaborative relationships among users and items. This existing CF paradigm can be enhanced by integrating encoded semantic representations. Typically, the preference score $p_{u,i}$ is calculated based on the interaction data, i.e., $p_{u,i}=f(u,i,\mathcal{A})$, where $\mathcal{A}$ represents the interaction data. Text-enhanced collaborative filtering builds upon this foundation by incorporating textual features $\mathbf{e}$ encoded by language models as supplementary representations. This integration aims to further improve the recommendation performance of the traditional ID-based collaborative filtering frameworks.
(3) $p_{u,i}=f(u,i,\mathcal{A},\mathbf{e}_{u},\mathbf{e}_{i}).$
3. Methodology
In this section, we first discuss how we gather textual profiles for users and items in recommender systems, which are essential for pre-training and evaluating our model. Next, we’ll dive into the specifics of EasyRec and its training approach. Lastly, we’ll introduce our method for diversifying user profiles, which improves the model’s ability to adapt to various situations.
3.1. Collaborative User and Item Profiling
In real-world recommenders, the only available information may be raw text data, such as product titles and categories, associated with the items. Privacy concerns often make it difficult to collect comprehensive user-side information. Furthermore, directly leveraging such textual information may overlook the crucial collaborative relationships needed for accurate user behavior modeling and preference understanding. To address these limitations, we propose to generate textual profiles by leveraging large language models (e.g., GPT, LLaMA series) (Ren et al., 2024a) to inject collaborative information into the textual profiles. This allows us to capture both the semantic and collaborative aspects of the items in a unified textual profile.
3.1.1. Item Profiling
Given the raw item information, such as title $h_{i}$, categories $c_{i}$, and description $d_{i}$ (e.g., book summary), we aim to generate a comprehensive item profile $\mathcal{P}_{i}$ that captures both the semantic and collaborative aspects. To reflect user-item interactions, we incorporate the textual information (e.g., posted reviews $r_{u,i}$) provided by the item’s corresponding users. Formally, the item profile generation process is:
(4) $\mathcal{P}_{i}=\text{LLM}(\mathcal{M}_{i},h_{i},c_{i},\{r_{u,i}\})\,\,\text{or}\,\,\text{LLM}(\mathcal{M}_{i},h_{i},d_{i}),$
Here, $\mathcal{M}_{i}$ represents the generation instruction, while $h_{i}$ and $d_{i}$ serve as input. By leveraging large language models (LLMs), we can generate a concise yet informative item profile $\mathcal{P}_{i}$.
3.1.2. User Profiling
In practical scenarios, privacy concerns often limit the feasibility of generating user profiles based on demographic information. Instead, we can profile users by considering their collaborative relationships, using the generated profile information from their interacted items. This approach allows the user profiles to effectively capture the collaborative signals that reflect their preferences. Formally, the user profile generation process is:
(5) $\mathcal{P}_{u}=\text{LLM}(\mathcal{M}_{u},\{h_{i},\mathcal{P}_{i},r_{u,i}|i\in\mathcal{N}_{u}\}).$
Here, $\mathcal{M}_{u}$ represents the instruction for generating the user profile using a large language model (LLM). We sample a set of interacted items $\mathcal{N}_{u}$ from the user’s purchase history. We then combine the user’s feedback $r_{u,i}$ with the pre-generated item profiles $\mathcal{P}_{i}$ to create the user’s text description $\mathcal{P}_{u}$, which captures their preferences.
3.1.3. Advantages of Collaborative Profiling
Our collaborative user and item profiling framework offers two key advantages for real-world recommendation scenarios which are elaborated as:
  • Collaborative Information Preserved. Our collaborative profiling approach goes beyond just the original textual content, also capturing the semantics of user/item characteristics and their interaction patterns. By encoding these rich profiles into a shared feature space using a recommendation-oriented language model, the resulting embeddings of interacted users and items are brought closer together. This enables the recommenders to better identify relevant matches, even for ”zero-shot” users and items (those without prior interactions) which are ubiquitous in real-world scenarios. The system can leverage the collaborative signals encoded within the text-based profiles to make better recommendations, bridging the gap for these cold-start cases.
  • Fast Adaptation to Dynamic Scenarios. Our user and item profiling approach, powered by robust language models, enables the recommender system to effectively handle the time-evolving nature of user preferences and interaction patterns. The key advantage is that simple updates to the textual user profiles can seamlessly reflect shifts in user interests and behaviors. This flexibility and responsiveness makes our approach well-suited for deploying recommender systems in dynamic environments where user interests can evolve over time.
3.2. Profile Embedder with Collaborative LM
So far, we have generated rich textual profiles for users and items, moving beyond conventional ID-based embeddings. However, directly encoding these textual profiles into latent embeddings for making recommendations may have two key limitations:
  • Capturing Recommendation-Specific Semantics. The text-based embeddings, while expressive, may not be optimized for the specific semantics and relationships most relevant to the recommendation task. For example, consider two item profiles: (1) “This user is passionate about advanced AI techniques, focusing on deep learning and AI research works”. (2) “With a passion for advanced AI development, this user delves into science fiction and AI-themed novels”. Though the profiles share textual similarity about AI, their target audiences differ - the first caters to AI scientists, the second to sci-fi readers. Directly encoding these profiles may overlook nuanced, recommendation-specific semantics. Refinement is needed to better align embeddings with the specific context and requirements of the recommendation system, beyond just textual similarity.
  • Overlooking High-Order Collaborative Signals. While textual profiles offer rich semantic information, relying solely on them may cause us to overlook valuable high-order collaborative patterns that emerge from complex user-item interactions (Wang et al., 2019; Xia et al., 2023). These higher-order signals, such as transitive associations and community-level preferences, can provide additional insights that complement the textual data for user preference learning in recommender systems.
To address these limitations, we propose a collaborative language modeling paradigm that seamlessly integrates the strengths of the semantic richness of the profiles and the valuable collaborative signals encoded from complex user-item interaction behaviors.
3.2.1. Bidirectional Transformer Encoder as Embedder
We leverage a multi-layer bidirectional Transformer encoder as the embedder backbone, considering two key benefits: 1) Efficient Encoding: The encoder-only architecture focuses solely on generating effective text representations, enabling faster inference in recommendation systems. 2) Flexible Adaptation: By building on pre-trained Transformer models, we can leverage transfer learning to optimize the embedder for specific recommendation tasks.
Let’s consider a user’s profile as a passage of $n$ words: $\mathcal{P}={w_{1},\ldots,w_{n}}$. We start by adding a special token [CLS] at the beginning of the word sequence. The tokenization layer then encodes the input sequence into initial embeddings, which serve as the input for the Transformer layers:
(6) $\{\mathbf{x}_{\texttt{[CLS]}}^{(0)},\ldots\mathbf{x}_{n}^{(0)}\}=\text{Tokenization}(\{w_{\texttt{[CLS]}},\ldots,w_{n}\}).$
Here, $x^{(0)}\in\mathbb{R}^{d}$ is the embedding retrieved from the embedding table corresponding to the tokens, and the $(0)$ superscript indicates that it is the input to the $(0)$-th layer of the language model. The tokenization process also adds positional embeddings to the initial embeddings. The language model then encodes a sequence of final embeddings (one for each token):
(7) $\{\mathbf{e}_{[CLS]},\ldots,\mathbf{e}_{n}\}=\text{Encoder}(\{\mathbf{x}_{\texttt{[CLS]}}^{(0)},\ldots\mathbf{x}_{n}^{(0)}\}),$
where Encoder($\cdot$) refers to the Transformer-based encoder-only language model. The key operation in the encoding process is the self-attention mechanism:
(8) $\text{Attention}(Q,K,V)=\text{softmax}({QK^{T}}/{\sqrt{d}})V$
[…]
Here, $X\in\mathbb{R}^{n\times d}$ represents the stack of token embeddings, and $W^{Q/K/V}$ are the parameter matrices that map these embeddings into queries, keys, and values. This self-attention mechanism allows each token to aggregate information from all other tokens, ensuring that each token is informed about the entire sequence. Finally, we select the first embedding $\mathbf{e}_{\texttt{[CLS]}}$, which corresponds to the [CLS] token, as the representative embedding for the entire profile. This embedding is then passed through a multi-layer perceptron to obtain the final encoded representation $e$, as mentioned in Eq.(1):
(10) $\mathbf{e}=\text{MLP}(\mathbf{e}_{\texttt{[CLS]}})=\text{LM}(\mathcal{P}).$
With these encoded text embeddings $\mathbf{e}$ for each user and item, we can predict the likelihood of interaction using cosine similarity and make recommendations as described in Eq.(2).
Figure 3. Contrastive tuning of collaborative language model empowers the model to learn rich representations that closely align the text-based semantic space with the globally collaborative signals from user behavior patterns.
3.2.2. Collaborative LM with Contrastive Learning
The motivation behind fine-tuning the collaborative language model using contrastive learning is to effectively capture and incorporate high-order collaborative signals into the recommendation model. Traditional recommenders using Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) optimize encoded embeddings with only one negative item per training sample. This approach limits the model’s ability to capture complex global user-item relationships.
The supervised contrastive loss provides a strong alternative to traditional methods. By treating interacted user-item pairs as positive views and non-interacted pairs as negatives, it brings related item embeddings closer together in the feature space. As shown in Figure 3, contrastive learning incorporates a batch of negatives, allowing for a comprehensive adjustment of the encoded feature space, which enables the model to capture high-order collaborative relationships. To evaluate this, we conduct experiments in Section 4.2.3 to assess the impact of different learning objectives. Specifically, the supervised contrastive loss for tuning the language model can be expressed as follows:
(11) $\mathcal{L}_{\text{con}}=-\frac{1}{|\mathcal{B}|}\sum_{(u,i_{\text{pos}},i_{\text{neg}})}\log\frac{\exp(\cos(\mathbf{e}_{u},\mathbf{e}_{i_{\text{pos}}})/\tau)}{\sum_{m\in\mathcal{N}_{\text{neg}}}\exp(\cos(\mathbf{e}_{u},\mathbf{e}_{m})/\tau)},$
Here, $\tau$ is a temperature hyperparameter that regulates the degree of contrastive learning, while $\mathcal{N}_{\text{neg}}$ denotes the set of in-batch negative items, which includes the negative item $i_{\text{neg}}$ and other items in the batch, excluding the positive item $i_{\text{pos}}$, with $|\mathcal{B}|$ depicts the batch size. Additionally, we follow prior work (Li et al., 2023; Hou et al., 2024) and incorporate an masked language modeling (MLM) loss $\mathcal{L}_{\text{mlm}}$. This technique randomly masks input tokens and trains the model to predict them, which helps stabilize the training process and enhance the model’s generalization ability. The final training objective is the combination of the supervised contrastive and the MLM loss:
(12) $\mathcal{L}=\mathcal{L}_{\text{con}}+\lambda\mathcal{L}_{\text{mlm}}.$
where $\lambda$ is a hyperparameter that controls the balance between the contrastive learning and the masked language modeling loss.
3.3. Augmentation with Profile Diversification
The goal of our profile diversification approach is to enhance the model’s ability to generalize to unseen users and items. Representing each user or item with a single profile inherently limits the diversity of the representations, which can negatively impact the model’s performance and generalization. To address this, we propose augmenting the existing user/item profiles to allow for multiple profiles per entity. These augmented profiles capture the same semantic meaning, such as the personalized interaction preferences of users or the varied characteristics of items. Our two specific augmentation methods introduce controlled variations in the profiles while preserving the core semantic meaning.
Inspired by self-instruction mechanisms (Wang et al., 2023; Xu et al., 2024), large language models (LLMs) can be leveraged to rephrase user or item profiles while preserving their underlying meaning. This allows generating multiple semantically similar yet distinctly worded profiles from a single input. Applying this iterative rephrasing process can create a diverse set of augmented profiles, substantially expanding the available training data. This data augmentation technique is particularly valuable when the original dataset is limited, as the LLM-generated profiles can improve model generalization and robustness.
By leveraging large language models for profile diversification, we can create a set of diverse profiles for each user $u$ and item $i$.
(13) $\{\mathcal{P}\}_{u}$ =$=\{\mathcal{P}_{u};\,\,\mathcal{P}_{u}^{1},\,\mathcal{P}_{u}^{2},\,\ldots,\,\mathcal{P}_{u}^{t}\},$
(14) $\{\mathcal{P}\}_{i}$ =$=\{\mathcal{P}_{i};\,\,\mathcal{P}_{i}^{1},\,\mathcal{P}_{i}^{2},\,\ldots,\,\mathcal{P}_{i}^{t}\}.$
Here, $\mathcal{P}_{u/i}$ represents the original profile, while $\mathcal{P}_{u/i}^{1-t}$ denotes the LLM-rephrased profiles, with $t$ indicating the number of diversification steps. During training, we randomly select a profile from the profile set for the user or item in each batch of user-item pairs.
4. Evaluation
This section evaluates the performance of the proposed EasyRec framework in addressing the following research questions (RQs):
  • RQ1: How effectively does the proposed EasyRec perform in matching unseen users and items (zero-shot) within text-based recommendation scenarios?
  • RQ2: How effectively does EasyRec integrate with and enhance various recommenders within text-based collaborative filtering scenarios, leveraging its capabilities as a language embedder?
  • RQ3: How effective is our proposed profile diversification mechanism for data augmentation in improving the performance of the recommendation language model?
  • RQ4: How well can our proposed text-based EasyRec paradigm adapt to accommodate changes in users’ dynamic preferences?
4.1. Experimental Settings
Table 1. Statistics of the datasets, with ”Avg. n” representing the average number of interactions per user. Datasets marked with underline are from different platforms.
Datasets #Users #Items #Inters. Avg. n Density
Train Data 124,732 67,455 802,869 6.44 -
[…]
Test Data 55,877 28,988 615,210 11.01 -
[…]
Table 2. The specifics of EasyRec’s model variants differ based on their model architecture and parameter sizes.
Model Layers Hidden size Heads Params
EasyRec-Small 6 768 12 82M
EasyRec-Base 12 768 12 125M
EasyRec-Large 24 1024 16 355M
Table 3. Text-based recommendation performance of various language models across different datasets. The best performance is indicated in bold, while the second-best is highlighted with underline. The script * denotes significance (p ¡ 0.05).
Data Sports Steam Yelp
Methods Recall@10 Recall@20 NDCG@10 NDCG@20 Recall@10 Recall@20 NDCG@10 NDCG@20 Recall@10 Recall@20 NDCG@10 NDCG@20
Proprietary Models
OpenAIv3-Small 0.0324 0.0444 0.0198 0.0230 0.0066 0.0119 0.0049 0.0068 0.0028 0.0056 0.0021 0.0031
OpenAIv3-Large 0.0300 0.0436 0.0180 0.0217 0.0070 0.0137 0.0049 0.0073 0.0029 0.0055 0.0023 0.0032
Base Size Models
BERT-Base 0.0015 0.0032 0.0008 0.0013 0.0015 0.0031 0.0011 0.0017 0.0009 0.0018 0.0006 0.0010
RoBERTa-Base 0.0121 0.0206 0.0065 0.0087 0.0041 0.0078 0.0031 0.0043 0.0018 0.0034 0.0014 0.0020
BART-Base 0.0088 0.0143 0.0048 0.0063 0.0033 0.0062 0.0024 0.0035 0.0014 0.0027 0.0012 0.0016
SimCSE-Base 0.0240 0.0341 0.0143 0.0170 0.0048 0.0091 0.0035 0.0050 0.0020 0.0039 0.0015 0.0022
BLaIR-Base 0.0251 0.0352 0.0145 0.0173 0.0041 0.0081 0.0030 0.0044 0.0025 0.0046 0.0019 0.0026
GTR-Base 0.0290 0.0417 0.0172 0.0206 0.0075 0.0136 0.0056 0.0077 0.0025 0.0052 0.0019 0.0029
BGE-Base 0.0315 0.0437 0.0194 0.0227 0.0094 0.0170 0.0069 0.0095 0.0029 0.0053 0.0022 0.0031
Large Size Models
BERT-Large 0.0011 0.0021 0.0006 0.0008 0.0019 0.0040 0.0014 0.0022 0.0008 0.0019 0.0007 0.0010
RoBERTa-Large 0.0039 0.0065 0.0021 0.0028 0.0027 0.0052 0.0021 0.0029 0.0012 0.0023 0.0010 0.0014
BART-Large 0.0114 0.0170 0.0064 0.0080 0.0036 0.0065 0.0026 0.0037 0.0017 0.0033 0.0014 0.0019
SimCSE-Large 0.0232 0.0328 0.0134 0.0160 0.0051 0.0095 0.0037 0.0052 0.0023 0.0044 0.0017 0.0025
BLaIR-Large 0.0227 0.0322 0.0133 0.0159 0.0057 0.0108 0.0041 0.0059 0.0027 0.0047 0.0019 0.0027
GTR-Large 0.0329 0.0446 0.0198 0.0229 0.0095 0.0168 0.0069 0.0094 0.0021 0.0042 0.0018 0.0025
BGE-Large 0.0324 0.0449 0.0196 0.0230 0.0089 0.0153 0.0066 0.0088 0.0025 0.0052 0.0020 0.0029
EasyRec Series
EasyRec-Small 0.0186 0.0286 0.0108 0.0135 0.0097 0.0174 0.0070 0.0097 0.0022 0.0046 0.0017 0.0026
EasyRec-Base 0.0360 0.0518 0.0210 0.0253 0.0114 0.0203 0.0081 0.0112 0.0034 0.0063 0.0026 0.0037
EasyRec-Large 0.0396* 0.0557* 0.0236* 0.0279* 0.0129* 0.0225* 0.0093* 0.0127* 0.0034* 0.0065* 0.0026* 0.0037*
Improve $\uparrow$ 20.36% $\uparrow$ 24.05% $\uparrow$ 19.19% $\uparrow$ 21.30% $\uparrow$ 35.79% $\uparrow$ 32.35% $\uparrow$ 39.13% $\uparrow$ 33.68% $\uparrow$ 17.24% $\uparrow$ 16.07% $\uparrow$ 13.04% $\uparrow$ 15.63%
Table 4. Comparison of EasyRec training with different optimization objectives (where ”Contrast” stands for contrastive).
Objective Sports Yelp
Recall@10 NDCG@10 Recall@10 NDCG@10
BPR Loss 0.0381 0.0226 0.0028 0.0021
Contrast Loss 0.0395 0.0236 0.0034 0.0026
4.1.1. Datasets
To assess our proposed model’s capability in encoding user/item textual profiles into embeddings for recommendation, we curated diverse datasets across various domains and platforms. A portion was used for training, while the remainder served as test sets for zero-shot evaluation. The dataset statistics are shown in Table 1. Due to the page limit, we place the detail of data resources are described in Appendix A.2.1
4.1.2. Evaluation Protocols
We employ two commonly used ranking-based evaluation metrics, Recall@$N$ and $N$, to assess performance in both text-based recommendation and collaborative filtering scenarios. Specifically, we compute these metrics for $N$ values of 10 and 20 (He et al., 2017, 2020). The evaluation is conducted using the all-rank protocol (He et al., 2020) with the predicted preference scores. In the context of text-based recommendation, particularly for datasets containing multiple LLM-diversified profiles (as discussed in Section 3.3), we calculate the metrics separately $t$ times based on different profile pairs $(\mathcal{P}_{u}^{1,\ldots,t},\mathcal{P}_{i}^{1,\ldots,t})$. Subsequently, we compute the mean value for each metric to obtain a comprehensive assessment. For the training datasets, we utilize the validation split for evaluation, while for the test datasets, we employ the test split.
4.2. Performance Comparision for Text-based Recommendation (RQ1)
We evaluate the performance of various language models (LMs) for zero-shot text-based recommendation on the unseen Sports, Steam, and Yelp datasets. This approach directly leverages the encoded embeddings derived from user/item profiles to make recommendations, without any additional training on the target datasets.
4.2.1. Baseline Methods and Settings
For our comparative evaluation, we included a diverse set of language models as text embedders: (i) General Language Models: BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), and BART (Lewis et al., 2019); (ii) Language Models for Dense Retrieval: SimCSE (Gao et al., 2021), GTR (Ni et al., 2021), and BGE (Xiao et al., 2023); (iii) Pre-trained Language Models for Recommendation: BLaIR (Hou et al., 2024). Additionally, we also compared against the state-of-the-art text embedding models provided by OpenAI. The detailed descriptions of these baseline models can be found in Appendix A.3. This selection of language models covers a broad spectrum, from general-purpose contextual encoders to specialized models tailored for tasks like dense retrieval and recommendation.
4.2.2. Result Analysis
The overall comparison of different models is presented in Table 3. This evaluation reveals several noteworthy observations, which are outlined below:
  • Superiority across Diverse Datasets. Our evaluation consistently shows that the EasyRec outperforms all other models across the three datasets spanning different platforms. This provides strong evidence for the effectiveness of the EasyRec. We attribute these improvements to two key factors: i) By injecting collaborative signals into the language models, we effectively optimized our EasyRec using supervised contrastive learning within the recommendation context. This approach allows the model to inherently encode user and item text embeddings that are well-suited for recommendation tasks. ii) By integrating a diverse array of datasets across multiple categories and utilizing data augmentation techniques to enrich the text descriptions for training, our EasyRec exhibits impressive generalization capabilities, enabling it to effectively handle unseen data.
  • Scaling Law Investigation of EasyRec Model. Our experiments revealed that as the size of the EasyRec model increases (from small to large), its performance consistently improves across all three datasets. This observation reflects a scaling law, where the model’s performance growth is directly correlated with its size. Furthermore, this finding effectively reinforces the validity of text-based recommendation systems. It also validates our approach to training the language model, which enables it to learn collaborative signals from a new perspective.
4.2.3. Impact of Training Objectives
Furthermore, we evaluate the impact of different training objectives on the language model’s learning process. To this end, we implemented EasyRec-Large training using BPR loss (i.e., one negative item per training sample) for comparison with the contrastive learning results. This approach allows us to directly assess how the choice of training objective influences model performance. As shown in Table 4, the performance of the model trained with contrastive learning generally outperforms that of the model trained with BPR loss. This outcome highlights the effectiveness of employing contrastive learning to better incorporate collaborative information into the language models, thereby enhancing their overall performance in recommendation tasks.
4.3. Performance of Text-enhanced CF (RQ2)
Table 5. Recommendation performance in text-enhanced collaborative filtering. The experiment was conducted on the Steam dataset with 5-runs to obtain the mean results.
Metric Recall NDCG
@5 @10 @20 @5 @10 @20
ID-based Methods
GCCF 0.0497 0.0826 0.1314 0.0555 0.0665 0.0830
LightGCN 0.0518 0.0851 0.1349 0.0575 0.0686 0.0854
Text-enhanced GCCF
BERT 0.0500 0.0822 0.1313 0.0556 0.0663 0.0829
RoBERTa 0.0517 0.0848 0.1351 0.0573 0.0684 0.0854
BART 0.0529 0.0874 0.1383 0.0585 0.0701 0.0874
SimCSE 0.0529 0.0877 0.1395 0.0588 0.0706 0.0881
BLaIR 0.0535 0.0880 0.1392 0.0593 0.0708 0.0882
GTR 0.0532 0.0873 0.1387 0.0592 0.0706 0.0880
BGE 0.0535 0.0875 0.1393 0.0591 0.0705 0.0881
EasyRec 0.0540 0.0881 0.1402 0.0597 0.0712 0.0888
Text-enhanced LightGCN
BERT 0.0518 0.0849 0.1347 0.0575 0.0684 0.0852
RoBERTa 0.0531 0.0867 0.1374 0.0587 0.0699 0.0870
BART 0.0540 0.0887 0.1407 0.0599 0.0715 0.0891
SimCSE 0.0541 0.0891 0.1417 0.0602 0.0719 0.0898
BLaIR 0.0551 0.0897 0.1418 0.0609 0.0724 0.0901
GTR 0.0542 0.0894 0.1417 0.0601 0.0719 0.0896
BGE 0.0547 0.0891 0.1407 0.0603 0.0718 0.0893
EasyRec 0.0554 0.0908 0.1430 0.0614 0.0732 0.0908
In addition to our investigation of zero-shot recommendation scenarios, we explore the potential of EasyRec as an enhancement when integrated with CF models. To assess the effectiveness of various LMs in CF, we employ two widely used ID-based methods as backbone models: GCCF (Chen et al., 2020) and LightGCN (He et al., 2020), which were chosen for their proven effectiveness and efficiency. Furthermore, we utilize the advanced model-agnostic text-enhanced framework RLMRec (Ren et al., 2024a) with contrastive alignment to conduct our investigation. We compare the large versions of both EasyRec and other open-source LMs. The key findings from the results in Table 5 are:
  • Compared to the backbone methods alone, the integration of the text-enhanced framework generally improves the performance for both GCCF and LightGCN. This observation highlights the significance of incorporating text modality into the recommendation paradigm, as it can enhance user and item profiles.
  • Among the various LMs tested, EasyRec consistently achieves the highest performance in the text-enhanced recommenders. This outcome not only illustrates the efficacy of EasyRec for recommendation tasks, but also emphasizes the advantages of incorporating collaborative information into language models.
4.4. Effectiveness of Profile Diversification (RQ3)
(a) Performance on Steam data
(b) Performance on Yelp data
Figure 4. Performance with respect to data size. ”Augmentation Count” indicates the number $t$ of diversified profiles.
In this section, we examine the impact of diversifying user and item profiles with large language models (LLMs) on model performance. As mentioned in Section 3.3, we perform LLM-based diversification three times on the original generated profiles. This process continuously increases the number of profiles in the training set. To investigate whether data augmentation positively affects model performance, we conduct experiments with three variants of the EasyRec under different numbers of diversified profiles. The results are shown in Figure 4, leading to the following key observations:
  • Effectiveness of Profile Diversification. The increase in the number of diversified profiles (from 0 to 3) enhances model performance, particularly for larger models. This finding underscores the effectiveness of our augmentation approach using LLMs for profile diversification, and emphasizes the significance of increasing training data for improved outcomes.
  • Scaling Relationship: The scaling experiments on both model size and data size reveal a crucial relationship that influences model performance. This demonstrates that our approach of training the language model with collaborative signals follows a scaling law, indicating that model performance benefits from both increased capacity and data volume. Such scaling laws are vital as they provide insights into how model capacity and data availability interact, guiding future research and development.
4.5. Model Fast Adaptation Case Study (RQ4)
Figure 5. Case study on handling user preference shift.
As mentioned in Section 3.1.3, a key advantage of EasyRec is its ability to empower recommender systems to efficiently adapt to shifts in user preferences and behavior dynamics over time. To evaluate this capability, we create two user profiles reflecting shifted preferences on the Amazon-Sport dataset and examine the recommended items from the EasyRec. As shown in Figure 5, the original user profile indicates that the user enjoys playing basketball. However, the user’s preference later transitions to a preference for swimming.
We visualize all the encoded embeddings using t-SNE (Van der Maaten and Hinton, 2008), which reveals a significant shift in the user embeddings within the feature space. Correspondingly, recommended items transition from basketball-related products to swimming gear, reflecting the user’s changing preferences. Notably, this adjustment is accomplished solely by modifying the user’s profile, without the need for further training of the model. This underscores the efficiency and flexibility of our approach in adapting to evolving user preferences.
5. Related Work
LMs-Powered Recommender Systems. Recent advancements in recommender systems increasingly incorporate textual modalities (Yuan et al., 2023), thus enhancing traditional approaches. The semantic representations encoded by pre-trained language models are essential features for improving recommender models, particularly in click-through rate prediction (Xi et al., 2023; Geng et al., 2024) and transferable sequence recommendations (Hou et al., 2022). These embeddings capture informative content relevant to recommendation tasks (Sheng et al., 2024). Some works also leverage text-based agents to enhance performance (Zhang et al., 2024a, b). A notable recent contribution is RLMRec (Ren et al., 2024a), a text-enhanced framework that improves ID-based recommenders using principles from information theory. However, many prior studies have relied on general text embeddings, such as BERT-based models (Ding et al., 2021; Yuan et al., 2023) or proprietary OpenAI embeddings (Ren et al., 2024a), rather than those specifically tailored for recommendation purposes. A recent work BLaIR (Hou et al., 2024) leverages the item metadata and interaction-level user feedback on this item for LMs training, yielding promising results in query-based item retrieval. In contrast, our approach assigns each user and item a collaboratively generated profile that reflects their preferences. EasyRec optimizes the learned correlations between user and item entities using CF signals, which not only demonstrates impressive zero-shot performance but also enhances text-augmented results.
Cross-Domain Recommendation. The fundamental concept behind cross-domain recommendation is to enhance recommendations in one domain by leveraging data from another domain, which is typically more abundant, to address data sparsity and improve personalization (Zang et al., 2022). For instance, the graph collaborative filtering network (Liu et al., 2020) has been introduced to aggregate both common and domain-specific user features using Graph Neural Networks (GNNs) that capture high-order user-item connections. Recent studies have further enriched cross-domain recommender systems through the integration of self-supervised learning techniques. For example, C2DSR (Cao et al., 2022) utilizes contrastive learning with both single-domain and cross-domain representations. CCDR (Xie et al., 2022) proposes intra-domain and inter-domain contrastive learning to enhance recommendation performance. Moreover, SITN (Sun et al., 2023) employs self-attention modules as sequence encoders to represent two user-specific sequences from the source and target domains, followed by contrastive learning between them. However, a significant limitation in current cross-domain research is its reliance on correlations (e.g., overlapping users) between source and target data, which constrains its applicability and generalizability. In contrast, text-based zero-shot learning does not has such constraint, allowing for effective transfer even across different datasets. Our proposed EasyRec takes a fundamental step forward in this line of research.
Graph Collaborative Filtering for Recommendation. Graph Neural Networks have recently emerged as a promising approach for recommendation systems, enabling the modeling of complex user-item interactions and capturing high-order dependencies in recommendation (Gao et al., 2023). These models are capable of learning representations of users and items by aggregating their interactions in a graph structure, such as PinSage (Ying et al., 2018), NGCF (Wang et al., 2019), and LightGCN (He et al., 2020). To enhance the representation capacity of GNNs against data sparsity in recommender systems, further studies realize the marriage between self-supervised learning and collaborative filtering with data augmentation. Examples include SGL (Wu et al., 2021), SimGCL (Yu et al., 2022), and HCCF (Xia et al., 2022). These developments showcase the potential of self-supervised graph learning in powering recommenders through graph augmentation based on node self-discrimination.
6. Conclusion
The EasyRec framework effectively integrates LMs to enhance recommendation tasks. Our new paradigm, which is both straightforward and effective, has consistently proven to excel across various scenarios, including text-based zero-shot recommendation and text-enhanced collaborative filtering. At the heart of EasyRec’s success lies an innovative methodology that combines collaborative language model tuning with the transformative capabilities of contrastive learning. This unique approach has empowered EasyRec to capture nuanced semantics and high-order collaborative signals—critical elements that have been instrumental in driving remarkable improvements in recommendation performance. Our extensive experiments, which span a diverse array of datasets, have consistently validated the superiority of EasyRec over existing language models. This robust and generalized performance underscores the framework’s remarkable capacity to adapt to dynamic user preferences, making it well-suited for real-world industry scenarios. Furthermore, the consistent improvements observed across different settings indicate that EasyRec is not only effective but also versatile in its application. Looking ahead, the potential for EasyRec to be seamlessly integrated with multi-modal information presents an enticing frontier for our future investigations.
Appendix A Appendix
User Profile Diversification Instruction You will assist me in revising a user’s profile while maintaining its original meaning. I will present you with the user’s initial profile.
Instructions: USER PROFILE: The original user profile.
Requirements: 1. Please provide the revised profile directly begin with "REVISED PROFILE: ". 2. The rephrased profile should minimize duplication with the original text while preserving its intended meaning. 3. The revised profile should exhibit varied sentence structures while faithfully conveying the original profile’s essence. Input Prompt USER PROFILE: This user is likely to enjoy items related to baking, entertaining, and colorful table settings. They appreciate convenience, efficiency, and practicality in kitchen appliances. Response REVISED PROFILE: An individual who finds pleasure in baking, hosting gatherings, and vibrant table arrangements. They value kitchen devices that offer convenience, efficiency, and practicality.
Figure 6. An example of large language models-based user profile diversification.
Item Profile Diversification Instruction You will assist me in revising a item’s profile while maintaining its original meaning. I will present you with the item’s initial profile.
Instructions: ITEM PROFILE: The original item profile.
Requirements: 1. Please provide the revised profile directly begin with "REVISED PROFILE: ". 2. The rephrased profile should minimize duplication with the original text while preserving its intended meaning. 3. The revised profile should exhibit varied sentence structures while faithfully conveying the original profile’s essence. Input Prompt ITEM PROFILE: The Innovee Lemon Squeezer is a high-quality stainless steel manual citrus press that comes with a lemon recipes ebook. Ideal for those who enjoy fresh lemon juice and recipes. Response REVISED PROFILE: The Innovee Lemon Squeezer is a stainless steel manual citrus press that includes a lemon recipes ebook, perfect for individuals who appreciate the taste of freshly squeezed lemon juice and love trying out new recipes.
Figure 7. An example of large language models-based item profile diversification.
A.1. Implementation and Training Details
We implemented our EasyRec and conducted all experiments using PyTorch (Paszke et al., 2019). For the transformer-based encoder backbone, we adopted the architecture of RoBERTa (Liu et al., 2019) and utilized its pre-trained parameters as initialization. We trained three versions of EasyRec with varying parameter sizes (small, base, and large), as detailed in Table 2. For the loss function, we set the hyperparameters $\tau$ to 0.05 and $\lambda$ to 0.1. The token masking ratio for masked language modeling is 0.15, and the learning rate is set to $5\times 10^{-5}$. We train the model for 25 epochs. For profile augmentation, we set the diversification time t for LLM-based methods to 3, resulting in 4 user profiles and 5 item profiles per dataset. During training, we evaluate the model every 1000 steps and use the validation interactions from each training dataset to select the optimal model parameters, employing the Recall@20 metric. Detailed implementation of our model is provided in our anonymous released code.
A.2. Datasets and User/Item Profiles
In this section, we provide detailed information on the datasets, as well as the processes for profile generation and diversification, including instructions, examples, and associated costs.
A.2.1. Details of Dataset
We utilize datasets from Amazon review data (Ni et al., 2019) across six categories to form the training data: Arts, Crafts and Sewing (Arts), Movies and TV (Movies), Video Games (Games), Home and Kitchen (Home), Electronics (Electronics), and Tools and Home Improvement (Tools). For the test datasets, we use one domain, Sports and Outdoors (Sports), from the Amazon review data, along with two cross-platform datasets: Steam and Yelp, for comprehensive evaluation. For the datasets from the Amazon platform, we first filter the data to include only those with a rating score greater than 3 and apply a 10-core filtering to densify the dataset. Subsequently, for each category, we split the interactions into training, validation, and test splits in a ratio of 8:1:1. In contrast, for the Steam and Yelp datasets, we directly use the data processed in previous work (Ren et al., 2024a), which maintains a split ratio of 3:1:1.
A.2.2. Details of Profile Generation
After data processing, each dataset contains a split of training interactions. We use these interactions to generate user and item profiles following the paradigm described in Section 3.1, as this requires user-item interaction information. For the Steam and Yelp datasets, we directly use the provided profiles, which adhere to the same generation protocol. It is important to note that, for each dataset, the profiles are generated exclusively based on the training interactions. This approach ensures that the validation and test interactions are reserved for evaluation purposes, preventing data leakage and allowing for a more accurate assessment of the model’s generalization performance on unseen recommendation data.
The instruction prompts for generation that we utilized reference previous research (Ren et al., 2024a) to conduct efficient user and item profiling. The profiling process adopts an item-to-user paradigm, where we first generate item profiles in parallel using multi-thread processing, followed by the parallel generation of user profiles that incorporate collaborative information. The large language model employed is GPT-3.5-Turbo from OpenAI.
Specifically, we process datasets from Amazon review data (Ni et al., 2019), which include the categories Arts, Movies, Games, Home, Electronics, Tools, and Sports. For item profile generation, each item $i$ includes a title $h_{i}$ and an original description $d_{i}$. We leverage these two pieces of information to generate the profile, as described on the left-hand side of Eq. 4. Next, for user profile generation, we uniformly sample a maximum of five interacted items from each user’s behavior history as collaborative information. We then arrange the input prompt for the large language model according to Eq. 5. All prompts and instructions for user and item profile generation are provided in the code for reference.
A.2.3. Details of Profile Diversification
As described in Section 3.3, we also conduct profile diversification using large language models (LLMs) to enhance the diversity of the training and test datasets, thereby improving and better evaluating the model’s generalization ability across different user and item profiles. For each user or item, we perform $t$ iterations of diversification starting from the initially generated profile. This means that we obtain the first diversified profile based on the original profile and then use this diversified profile for further diversification with the LLMs.
For reference, examples of user and item profile diversification are provided in Figure 6 and Figure 7, respectively. As illustrated in the case of user profile diversification, the profiles for the same user differ at the word level while still representing the same preferences. Such diversification significantly enhances the diversity and quality of the textual data while also increasing the overall dataset size.
Table 6. Costs for user and item profile generation and diversification, including three iterations of diversification.
Operation #Data #In. Tokens #Out. Tokens #Cost (\$)
Generation 7 169M 20M $\sim$114
Diversification 9 102M 30M $\sim$97
A.2.4. Cost of Generation and Diversification
We summarize the total number of tokens and the associated costs for utilizing the proprietary model for profile generation and diversification in Table 6. It is worth noting that the profiles for the Steam and Yelp datasets have already been generated; therefore, the number of profiled datasets and diversified datasets differs. As shown in the results, the total number of tokens required to process both profile generation and diversification is approximately 322 million, including both input and output tokens. This generally incurs a cost of around 200 dollars with GPT-3.5-Turbo API to process the entire dataset, making it an affordable option.
A.3. Details of Text-based Recommendation
Table 7. Details of compared language models.
Model Pre-trained Weights (From Hugging Face)
BERT-Base google-bert/bert-base-uncased
BERT-Large google-bert/bert-large-uncased
BART-Base facebook/bart-base
BART-Base facebook/bart-large
RoBERTa-Base FacebookAI/roberta-base
RoBERTa-Large FacebookAI/roberta-large
SimCSE-Base princeton-nlp/sup-simcse-roberta-base
SimCSE-Large princeton-nlp/sup-simcse-roberta-large
BLaIR-Base hyp1231/blair-roberta-base
BLaIR-Large hyp1231/blair-roberta-large
GTR-Base sentence-transformers/gtr-t5-base
GTR-Large sentence-transformers/gtr-t5-large
BGE-Base BAAI/bge-base-en-v1.5
BGE-Large BAAI/bge-large-en-v1.5
A.3.1. Baseline Models
In this section, we provide a detailed description of the language models compared in this work.
(i) General Language Models.
  • BERT (Devlin et al., 2018): A landmark transformer-based model renowned for strong language understanding through bidirectional training. We use the pooled BERT output as the text embedding.
  • RoBERTa (Liu et al., 2019): An optimized BERT that employs dynamic masking and larger datasets. We use the final [CLS] token embedding.
  • BART (Lewis et al., 2019): A denoising autoencoder transformer trained on corrupted text reconstruction. We apply mean pooling on the last hidden state for the text embedding.
(ii) Language Models for Dense Retrieval.
  • SimCSE (Gao et al., 2021): A framework that leverages contrastive learning to generate high-quality sentence embeddings, enhancing the model’s ability to discern semantic similarity between sentences.
  • GTR (Ni et al., 2021): GTR is a generalizable T5-based dense retriever that improves retrieval tasks across various domains by overcoming limitations of traditional dual encoders.
  • BGE (Xiao et al., 2023): BGE is a state-of-the-art family of well-trained models for general text embedding models. We utilize the English version of this model for evaluation.
(iii) Langauge Models for Recommendation.
  • BLaIR (Hou et al., 2024): A series of sentence embedding models for recommendation. BLaIR learns interaction-level correlations between item metadata and user feedback on this item, improving query-based item retrieval and also recommendation.
The pre-trained weights utilized for each baseline language model are listed in Table 7. For proprietary embedding models named OpenAIv3, we use the latest text-embedding-3-small and text-embedding-3-large from OpenAI as baselines.
A.3.2. Detail of Baseline Settings
Given that the experiment was conducted in a zero-shot setting, where the test data remained unseen during training for both our EasyRec and the other language model baselines, we directly utilized the original released parameters of these baselines from Hugging Face for initialization and comparison. Besides, for BGE, we added the recommended retrieval instruction in front of the user profiles, following the provided guidelines from the open-source code, as we found it offered better performance in zero-shot recommendations. Additionally, we applied normalization to the outputs from the language models to derive text embeddings and computed cosine similarity using the dot product of the user and item embeddings.